The perimeter of the cover of an address book is 14 inches, and the diagonal measures 5 inches. What are the length and width of the cover?
Answer by RAY100(1637)

perimeter = 2l +2w =14
.
l+w=7,,,or w=7-l
.
.
diagonal ^2 = l^2 + w^2= 5^2 =25,,,,,subst
.
l^2 +(7-l)^2 =25
.
l^2 +49 -14l +l^2 = 25
.
2l^2 -14l +24 =0,,,,,divide by 2
.
l^2 -7l +12 = 0,,,,,factor
.
(l-4)(l-3) = 0
.
l=4,3,,,,,and w= 3,4
.
or conventionally (w smaller) w=3,l=4
.
check ,,,,dia = 5 = sqrt{ 3^2 +4^2} ,,,,,ok
.
per = 2(3+4)= 14,,,,,,ok
,
.